if the sum of the squares of the distances of a point from 3 co-ordinate axis is 36 then find the distance from origin

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
The sum of squares of distance from the point to each coordinate axis = 36.
Distance from a point (x,y,z) to x-axis = x.
Distance from a point (x,y,z) to y-axis = y.
Distance from a point (x,y,z) to z-axis = z.
We have that x2 + y2 + z2 = 36.
Distance from origin to the point (x,y,z) = sqrt( (x-0)2 + (y-0)2 + (z-0)2 ) = sqrt (x2 + y2 + z2) = 6.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)